### v2

# Competition between group interactions and nonlinearity in voter dynamics on hypergraphs

Abstract
Social dynamics are often driven by both pairwise (i.e., dyadic) relationships and higher-order (i.e., polyadic) group relationships, which one can describe using hypergraphs. To gain insight into the impact of polyadic relationships on dynamical processes on networks, we formulate and study a polyadic voter process, which we call the group-driven voter model (GVM), that incorporates the effect of group interactions by nonlinear interactions that are subject to a group (i.e., hyperedge) constraint. By examining the competition between nonlinearity and group sizes, we show that the GVM achieves consensus faster than standard voter-model dynamics, with an optimal minimizing exit time. We substantiate this finding by using mean-field theory on annealed uniform hypergraphs with $N$ nodes, for which the exit time scales as ${\cal A}\ln N$, where the prefactor ${\cal A}$ depends both on the nonlinearity and on group-constraint factors. Our results reveal how competition between group interactions and nonlinearity shapes GVM dynamics. We thereby highlight the importance of such competing effects in complex systems with polyadic interactions.
Introduction—Individuals in society interact both in pairs and through various types of social groups (including families, clubs, and work colleagues) Olmsted (1962); Wheelan (1994). Group (i.e., “polyadic”) interactions often are not merely structural units of a network; they also constitute functional units that drive dynamics through nonlinear effects Tyson (1998); Brown and Pehrson (2019). Consequently, the traditional network framework — which employs graphs and thus encodes group interactions as collections of pairwise (i.e., dyadic) interactions Bick et al. (2023) — has a fundamental limitation. To explicitly capture group interactions, one can employ “higher-order” (i.e., polyadic) network frameworks Bick et al. (2023); Battiston et al. (2020); Bianconi (2021); Boccaletti et al. (2023). There has been much recent work on dynamical processes on polyadic networks Majhi et al. (2022); Iacopini et al. (2019); Skardal and Arenas (2019); St-Onge et al. (2021, 2022); de Arruda et al. (2023); Carletti et al. (2023); Cencetti et al. (2023); Kim et al. (2023); Burgio et al. (2024); Kim and Goh (2024); Civilini et al. (2024) on a variety of systems, including opinion dynamics Sahasrabuddhe et al. (2021); Noonan and Lambiotte (2021); Schawe and Hernández (2022); Hickok et al. (2022); Horstmeyer and Kuehn (2020); Papanikolaou et al. (2022). However, researchers have not obtained a generic understanding of the impact of group interactions on dynamical processes.
To gain insight into the impact of group interactions on opinion dynamics, we formulate and analyze a polyadic voter model. Voter models (VMs) Liggett (1999) are both among the simplest models of social dynamics Castellano et al. (2009a) and among the best-understood theoretical models of collective behavior of complex systems Krapivsky et al. (2010); Dornic et al. (2001). One can interpret the update rules of a VM  [^1] There are several variants of “the” VM, depending on choices such as whether one selects nodes or edges at random, that have different qualitative dynamics Redner (2019). in terms of choosing between binary choices, which we denote by $\sigma=\mathbf{0}$ and $\sigma=\mathbf{1}$. We use the terms “opinion” and “state” interchangeably for the variable $\sigma$. At each time step, a uniformly random node adopts the opinion of a uniformly random neighbor. (Henceforth, we use the term “random” as a shorthand description for uniformly at random.) VMs have been studied for more than half of a century Holley and Liggett (1975), and they have been studied actively on traditional networks (i.e., graphs) for more than two decades Castellano et al. (2003); Suchecki et al. (2005); Sood and Redner (2005); Vazquez et al. (2008); Sood et al. (2008); Vazquez and Eguíluz (2008); Masuda (2014). VMs have also been extended in a variety of ways Redner (2019). However, few existing studies account explicitly for group interactions Horstmeyer and Kuehn (2020); Papanikolaou et al. (2022).
The framework of polyadic networks can help fill this gap by providing explicit structural models, such as hypergraphs and simplicial complexes Bick et al. (2023), to deal with group interactions. To incorporate group interactions into opinion dynamics, we use hypergraphs and generalize VM dynamics. To initiate our generalization, we first reformulate a traditional dyadic VM update rule by focusing on the role of edges. At each time step, a random node $i$ chooses one of its edges (i.e., links) at random, and it flips its state $\sigma_{i}$ to the state $\sigma_{j}$ of the adjacent node $j$ that is attached to the chosen edge if the node states are different. In a dyadic network, each of these edges of a node $i$ is attached to exactly 1 other node.
[…]
Group-driven voter model—In a hypergraph, a node $i$ can be adjacent to more than one other node via a hyperedge. Each node in a hyperedge with cardinality (i.e., “size”) $s$ is adjacent to $s-1$ nodes. This multiplicity leads to a broad spectrum of possibilities for dynamical processes on hypergraphs. To investigate these possibilities, we study a group-driven voter model (GVM). At each time step, a random node $i$ considers adopting an opinion from one of its incident hyperedge $h$, which we choose randomly. During the adoption process, node $i$ makes $q$ observations of states (i.e., opinions) of random nodes of hyperedge $h$. One can either allow Castellano et al. (2009b) or disallow Centola and Macy (2007) duplicate choices of the same neighbor. If the $q$ observed node states $\𝑖\{\sigma_{j_{1}},\sigma_{j_{2}},\ldots,\sigma_{j_{q}}\,|\,j_{p}\in h\backslash\{i\}\}$ are unanimous and different from its own state $\sigma_{i}$, then node $i$ flips its state (see Fig. 1) to match the observed state.
The GVM has two independent parameters: $q$ and $s$. The parameter $q$ accounts for nonlinear interactions Nowak et al. (1990), which are absent in standard VMs but have been considered in nonlinear variants of voter models Lambiotte and Redner (2007); Castellano et al. (2009b); Volovik and Redner (2012); Kureh and Porter (2020); Ramirez et al. (2024). The parameter $s$ accounts for the effect of polyadic interactions. The GVM incorporates social reinforcement Bond (2014); Centola and Macy (2007); Ruan et al. (2015); Juul and Porter (2018) via group interactions Neuhäuser et al. (2020); St-Onge et al. (2021), suggesting an explicit group-based origin of nonlinearity, which has been introduced in an ad hoc way in various dyadic variants of VMs Jedrzejewski and Sznajd-Weron (2019), including a vacillating voter model Lambiotte and Redner (2007), a $q$-voter model Castellano et al. (2009b), a confident voter model Volovik and Redner (2012), and a nonlinear voter model Ramirez et al. (2024).
When $q=s-1$ and duplicate choices are disallowed, the GVM captures the strongest group interactions, as it requires that all of the $s-1$ nodes’ states of a selected hyperedge are unanimous and different from node $i$’s state for node $i$ to flip its state. This requirement amounts to a “simplicial rule”, which was used in Refs. Iacopini et al. (2019); Robiglio et al. (2024) to refer to polyadic interactions that require unanimity of states. We thus refer to this variant as a “simplicial GVM”. When $q=1$ for all values of $s$, the GVM essentially reduces to a standard dyadic VM; it no longer experiences the effects of polyadic interactions. Additionally, the GVM on dyadic networks ($s=2$) reduces to the standard VM for all values of $q$. When $s=N$, where $N$ is the number of nodes, the GVM reduces to the noiseless $q$-voter model Castellano et al. (2009b); Ramirez et al. (2024) on a fully-connected dyadic network. However, for networks that are not complete, the correspondence is not exact due to the explicit group constraint.
Figure 2: The exit time $\tau$ for the simplicial GVM on annealed hypergraphs with hyperedge sizes that we obtain from (a) a geometric distribution with different mean hyperedge sizes $\langle s\rangle$ and (b) a power-law distribution with different power-law exponents $\alpha$. The symbols and shaded areas indicate the means and standard deviations, respectively, of $10^{3}$ independent Monte-Carlo simulations of the simplicial GVM with $N=10^{5}$ nodes, and the curves indicate analytical results from a recursion relation (solid curves) and a leading-order approximate solution (dotted curves). (See the Supplemental Material (SM) for the associated equations SM .) Both situations exhibit a notion of optimality: With increasing $\langle s\rangle$ or decreasing $\alpha$, the exit time $\tau$ first decreases but eventually increases, achieving a minimum in the middle.
To clearly observe the effect of groups, we consider the simplicial GVM. A basic property of voter dynamics is the exit time, which is the time that it takes to reach consensus (of either state). The exit time depends on the initial node states. Let $\tau$ denote the exit time from a balanced initial condition, in which the same number of nodes are in each state. In Fig. 2, we show the exit time $\tau$ for the simplicial GVM on “annealed” hypergraphs, in which the nodes of a hyperedge are determined uniformly at random in each time step Boguñá et al. (2009), with two different hyperedge-size distributions $P(s)$ — a geometric distribution $P(s)=[(\langle s\rangle-2)/(\langle s\rangle-1)]^{s-2}/(\langle s\rangle-1)$ for $s\geq 2$ and a power-law distribution $P(s)=s^{-\alpha}/\sum_{\ell=2}^{\infty}\ell^{-\alpha}$ for $s\geq 2$ — that are inspired by empirical data sets Patania et al. (2017); Benson et al. (2018); Roh and Goh (2023). We compute the exit time $\tau$ as a function of the mean hyperedge size $\langle s\rangle$ (for the geometric distribution) and the power-law exponent $\alpha$ (for the power-law distribution) using Monte-Carlo (MC) simulations of the simplicial GVM on hypergraphs with $N=10^{5}$ nodes  [^2] In the SM SM , we give the algorithmic details of our Monte-Carlo simulations. Associated code is available at https://github.com/JihyeKim2024/GVM.. In both cases, $\tau$ behaves nonmonotonically (see Fig. 2), so there are optimal values of $\tau$. As groups of three or more nodes begin to appear (i.e., $\langle s\rangle\gtrapprox 2$ or $3\lessapprox\alpha<\infty$), consensus accelerates (i.e., $\tau$ becomes smaller). However, when group sizes are too large (i.e., $\langle s\rangle\gg 2$ or $\alpha\lessapprox 3$), consensus decelerates. Therefore, there is an “optimal” level of group interactions that leads to the fastest consensus (i.e., the smallest $\tau$).
To gain theoretical insight into the origin of this optimality, we henceforth analyze the GVM with duplicate choices allowed on annealed $s$-uniform hypergraphs with $N$ nodes. We use this setting because it allows us to derive a series of concrete, informative analytical results. For simplicity, we treat all nodes as equivalent; at each time step, we select nodes of a hyperedge uniformly at random with replacement. In an $s$-uniform hypergraph, each hyperedge has the same size (i.e., the same number of nodes) $s$. We consider the competition between our two independent parameters, $q$ and $s$, in the opinion dynamics. In this case, $q\geq s$ is also possible because we allow duplicate selections of the same neighboring node.
Mean-field theory—To theoretically understand the GVM dynamics, we use mean-field theory Sood et al. (2008); Ramirez et al. (2024). A key variable is the density $\rho(t)$, which is the fraction of nodes of a hypergraph in state $\mathbf{1}$ at time $t$. In a time step, $\rho(t)$ can increase or decrease by $\delta\rho=1/N$. One can account for this change with the transition probability $R(\rho)\equiv P(\rho\rightarrow\rho+\delta\rho)$ that the number of nodes in state $\mathbf{1}$ increases by $1$ in a time step and the transition probability $L(\rho)\equiv P(\rho\rightarrow\rho-\delta\rho)$ that it decreases by $1$ in a time step. The probability of no change in $\rho$ in one time step is $1-R(\rho)-L(\rho)$. The rate equation for $\rho(t)$ is
$\dfrac{d\rho}{dt}=R(\rho)-L(\rho)\equiv v(\rho)\,,$ (1)
where $v(\rho)$ is the drift function.
For an annealed $s$-uniform hypergraph, one can write SM
$R(\rho)$ $=(1-\rho)\sum_{n=0}^{s-1}{s-1\choose n}\rho^{n}(1-\rho)^{s-1-n}\left(\frac{n}{s-1}\right)^{q}$
$=\left.\frac{(1-\rho)}{(s-1)^{q}}\left(\frac{d}{dr}\right)^{q}\left[(1-\rho+\rho e^{r})^{s-1}\right]\right|_{r=0}\,,$
$L(\rho)$ $=\rho\sum_{n=0}^{s-1}{s-1\choose n}\rho^{n}(1-\rho)^{s-1-n}\left(1-\frac{n}{s-1}\right)^{q}$
$=\left.\frac{\rho}{(s-1)^{q}}\left(\frac{d}{dr}\right)^{q}\left[(\rho+e^{r}-\rho e^{r})^{s-1}\right]\right|_{r=0}\,.$ (2)
In this mean-field approximation, the probability that a size-$s$ hyperedge has $n$ nodes in state $\mathbf{1}$ at time $t$ is (${s-1\choose n}\rho^{n}(1-\rho)^{s-1-n}$.
The drift function $v(\rho)$ gives many useful insights about GVM dynamics. When $q=1$ (i.e., for the standard VM), $v(\rho)=0$ for all $\rho$ because $R(\rho)=L(\rho)=\rho(1-\rho)$ Sood et al. (2008). In this case, stochastic fluctuations enable finite-size systems to reach consensus. For the generic GVM (i.e., when $q\geq 2$), the drift function $v(\rho)$ is no longer identically 00. We show in the Supplemental Material (SM) SM that Eq. (1) has three equilibrium points: $\rho=0$, $\rho=1$, and $\rho=1/2$. The equilibrium points $\rho=0$ and $\rho=1$ are stable and correspond to consensus states with opinions $\mathbf{0}$ and $\mathbf{1}$, respectively. Apart from finite-size fluctuations, the system eventually reaches the $\rho=0$ consensus equilibrium whenever $\rho<1/2$ because $v(\rho)<0$. For $\rho>1/2$, the system eventually reaches the consensus equilibrium $\rho=1$. The unstable equilibrium point $\rho=1/2$ has an equal mixture of the opinions $\mathbf{0}$ and $\mathbf{1}$. Drift towards a stable equilibrium point depends on the values of $q$ and $s$, which thereby play crucial roles in the GVM dynamics. The drift function $v(\rho)$ of the GVM for $s=N$ reduces to that of the $q$-voter model on a fully-connected dyadic network SM .
Sigmoidal exit probability—Another key property of voter dynamics is the exit probability $\Phi(\rho)$, which is the probability to reach the opinion-$\mathbf{1}$ consensus state from the initial density $\rho$. From the preceding argument, we expect that the exit probability for the generic GVM (i.e., for any $q\geq 2$) changes in a sigmoidal manner near $\rho=1/2$, with convergence to a step function in the thermodynamic limit $N\to\infty$, as has also been observed in numerical simulations of the $q$-voter model Castellano et al. (2009b). To confirm this expectation and elucidate the group effect, we calculate $\Phi(\rho)$ explicitly for large but finite $N$. Following Sood et al. (2008), we set up the recursion relation
$\Phi(\rho)$ $=R(\rho)\Phi(\rho+\delta\rho)+L(\rho)\Phi(\rho-\delta\rho)$
$\quad\,+\left[1-R(\rho)-L(\rho)\right]\Phi(\rho)$ (3)
and Taylor-expand it in $\delta\rho=1/N$ to second order to obtain a backward Kolmogorov equation
$v(\rho)\dfrac{\partial{\Phi(\rho)}}{\partial{\rho}}+D(\rho)\dfrac{\partial^{2}{\Phi(\rho)}}{\partial{\rho^{2}}}=0\,,$ (4)
with a diffusion function $D(\rho)\equiv[R(\rho)+L(\rho)]/(2N)$ and boundary conditions $\Phi(0)=0$ and $\Phi(1)=1$. By symmetry, $\Phi(1-\rho)=1-\Phi(\rho)$.
Figure 3: The exit probability $\Phi(\rho)$ when the nonlinearity strength is $q=2$ and the group size is $s=3$. The solid curves are from the function in Eq. (5), and the markers are means of $10^{4}$ independent MC simulations of the GVM on annealed $3$-uniform hypergraphs. As $N$ increases, the sigmoid $\Phi$ converges to a step function. In the inset, we show $\Phi(\rho)$ for a fixed system size $N$ and different group sizes $s$. Convergence to a step function is slower for $s=3$ than for $s=N$.
To illustrate the effect of the nonlinearity, we compare the two simplest cases: $q=1$ (i.e., the standard VM) and $q=2$ (our GVM). When $q=1$, it is known that $\Phi(\rho)=\rho$ Sood et al. (2008), as one can view the dynamics as a diffusion process (i.e., $v(\rho)=0$). For $q=2$, we solve Eq. (4) explicitly to obtain SM
$\Phi(\rho)=\dfrac{1}{2}+\dfrac{\mathrm{erf}\left(\sqrt{\dfrac{2N(s-2)}{s}}\left(\rho-\dfrac{1}{2}\right)\right)}{2\,\mathrm{erf}\left(\sqrt{\dfrac{N(s-2)}{2s}}\right)}\,,$ (5)
where $\mathrm{erf}(\cdot)$ is the error function. In Fig. 3, we plot $\Phi(\rho)$ when $q=2$ and $s=3$. This expression agrees with the results of our MC simulations. This explicit closed-form confirmation demonstrates that the “width” $\Delta$ of the sigmoidal change across $\rho=1/2$ scales as $\Delta\sim 1/\sqrt{N(s-2)/s}$, illustrating both the finite-size effects (i.e., the dependence on $N$) and the group effect (i.e., the dependence on $s$). In particular, we see that convergence to a step function “slows down” for smaller group sizes $s$ when $q=2$. See the inset of Fig. 3.
Logarithmic scaling of the exit time $\tau$ with hypergraph size $N$—Let $T(\rho)$ denote the exit time for a general initial density $\rho$ of nodes in state $\mathbf{1}$. Therefore, $\tau=T(\rho=1/2)$. Following a similar procedure as in our derivation of Eq. (4) for the exit probability, we set up a recursion relation for the exit time $T(\rho)$. This yields the backward Kolmogorov equation SM
$v(\rho)\dfrac{\partial{T(\rho)}}{\partial{\rho}}+D(\rho)\dfrac{\partial^{2}{T(\rho)}}{\partial{\rho^{2}}}=-1\,.$ (6)
For the standard VM (i.e., for $q=1$), the drift term vanishes and we solve Eq. (6) and obtain $\tau=N\ln 2$  Sood et al. (2008); SM . However, for the GVM (i.e., for $q\geq 2$), it is typically not possible to solve Eq. (6) analytically. Nevertheless, one can numerically solve the recursion relation for $T(\rho)$ that is analogous to Eq. (3). See Eq. (S21) in the SM SM .
To proceed further analytically, we approximate Eq. (6) by neglecting the diffusion term. We can do this because $D(\rho)/v(\rho)\sim 1/N$ as $N\rightarrow\infty$. We then integrate the resulting equation to obtain the approximate exit time
[…]
We have shifted the initial density by $1/\sqrt{N}$ from $1/2$ to exploit stochasticity and thereby avoid being trapped at the unstable equilibrium point. Under this approximation, we obtain to leading order in the hypergraph size $N$ that $\tau(N;s,q)\sim{\cal A}(s,q)\ln N$, where the prefactor ${\cal A}(s,q)$ depends on $s$ and $q$ for general $q\geq 2$ and $s\geq 3$. (See Eq. (S33) in the SM SM .) One can attribute the logarithmic scaling of the exit time for the generic GVM to the fact that $v(\rho)=0$ has three simple roots in $[0,1]$. The prefactor ${\cal A}(s,q)$ diverges for $s=2$, as $\tau$ satisfies diffusive scaling $\tau\sim{\cal O}(N)$ for dyadic networks.
Figure 4: Logarithmic scaling of the exit time $\tau(s,q,N)$ for the GVM on the hypergraph size $N$ for (a) nonlinearity strength $q=2$ and (b) nonlinearity strength $q=5$. The markers are means of $10^{6}$ (when $N\leq 10^{4}$) or $10^{3}$ (when $N\geq 10^{5}$) independent MC simulations of the GVM on annealed $s$-uniform hypergraphs. We obtain the solid curves from Eq. (S21) in the SM SM , and we obtain the dashed curves from the leading-order solutions in Eqs. (8, 9). The dashed and solid green curves ($s=7$) in (a) almost overlap.
It is insightful to show some explicit approximate expressions for $\tau$ as examples. The leading-order expression of $\tau$ for $q=2$ and $q=5$ are SM
$\tau(N,s;q=2)$ $\sim\dfrac{2(s-1)}{(s-2)}\ln N\,,$ (8)
$\tau(N,s;q=5)$ $\sim\dfrac{(s-1)^{4}\dfrac{(3s-4)(s+1)}{\left(s^{2}+3s-8\right)}}{s(s-2)(s^{2}-2s+2)}\ln N\,.$ (9)
From Fig. 4, we see that the analytically-obtained logarithmic scaling of $\tau$ successfully explains the MC simulation results. Figure 4 also reveals that the group effect can manifest distinctively for different nonlinearity strengths $q$. When $q=2$, reaching consensus takes the longest time for the smallest group size $s=3$ [see Fig. 4(a)]. By contrast, when $q=5$, the longest consensus time occurs for the largest group size $s=N$ [see Fig. 4(b)].
Figure 5: (a) The dependence on the nonlinearity strength $q$ of the exit time $\tau$ for the GVM for different values of the group size $s$ for hypergraphs with $N=10^{4}$ nodes. The markers are means of $10^{4}$ independent MC simulations of the GVM on annealed $s$-uniform hypergraphs, and the analytical curves are from numerical solutions of the recursion relation (S21) in the SM SM . When $s=N$, the exit time $\tau$ grows exponentially quickly with $q$, making it problematic to depict it along with other cases. In the inset, we show the curve for $s=N$ in an extended $q$ range. (b) The heat map for $\tau(s,q,N)$ that we obtain from numerical solutions of the recursion relation (S21). For a given $s$, the cell with the red border has the optimal $\tau$.
Optimality in the exit time $\tau$—To further examine the interplay between the nonlinearity and group effects on the exit time $\tau$, we investigate how varying the nonlinearity strength $q$ affects the GVM dynamics for specified values of the group size $s$ and hypergraph size $N$. In the absence of the group constraint (i.e., $s=N$), the leading-order expression of the exit time $\tau$ is
$\tau(q;N,s=N)\sim\left(1+\frac{2^{q-2}}{q-1}\right)\ln N\,.$ (10)
This expression also applies to the $q$-voter model on a complete dyadic graph. Because $\tau$ increases with $q$, a stronger nonlinearity decelerates consensus. [See Fig. 5(a) and its inset.] By contrast, with the most-constraining groups (i.e., $s=3$), the exit time $\tau$ decreases monotonically with $q$. To leading order, $\tau(q;N,s=3)\sim\frac{2^{q}}{2^{q-1}-1}\ln N$. A stonger nonlinearity accelerates consensus. [See the red curve in Fig. 5(a).]
There is a nontrivial tradeoff between these extreme situations. As we can see in Fig. 5(a), for a given hyperedge size $s$, there is an optimal nonlinearity strength $q^{*}$ with the minimum exit time $\tau$. We systematically investigate the tradeoff for many values of $s$ and $q$ [see Fig. 5(b)]. These computations reveal the global landscape and optimality of GVM dynamics.
We now explain why we observe optimality. From the inset of Fig. 5(a), we see that considering the opinions of exactly $2$ neighbors is the most efficient way to achieve consensus. Increasing $q$ in Eq. (10) with $s=N$ reduces the probability that neighbors have unanimous opinions, which in turn decreases the drift function $v(\rho)$ and decelerates the approach to consensus. Because $q$ grows logarithmically with $N$ [i.e., $q\sim{\cal O}(\ln N)$], the logarithmic scaling of $\tau$ in Eq. (10) eventually becomes a linear scaling $\tau\sim{\cal O}(N)$, which is comparable to the diffusive scaling for $q=1$. However, the probability that a node consults the same neighbor twice (instead of consulting 2 different neighbors) for $q=2$ is $1/(s-1)$, which is not negligible for small $s$. In this situation, a node only consults the opinion of $1$ neighbor, so it again effectively follows diffusive dynamics. More generally, the probability of diffusive dynamics from consulting just $1$ neighbor increases with decreasing $q$. Therefore, there is a “sweet spot” $q^{*}$ that minimizes $\tau$ between the two diffusive-dynamics maxima. That is, $2<q^{*}<{\cal O}(\ln N)$. The case $s=3$ is a notable exception. When $s=3$, the maximum number of different neighbors is $2$, so $\tau$ decreases indefinitely (although slowly) as $q$ increases.
When $q$ is fixed, increasing $s$ towards $N$ reduces the probability of unanimity, as the number of distinct neighbors that are chosen increases, and decreasing $s$ towards $2$ increases the probability of consulting just $1$ neighbor. Both situations lead to an increase in the exit time $\tau$. Therefore, there exists an optimal $\tau$. An equivalent explanation of the presence of optimality in the simplicial GVM (with $q=s-1$ and duplicate choices disallowed) in Fig. 2 is as a competition between diffusive dynamics from dyadic edges (which dominates as $\langle s\rangle\downarrow 2$ and $\alpha\to\infty$) and the small probability of unanimity in large hyperedges (which dominates as $\langle s\rangle\to\infty$ and $\alpha\downarrow 2$). The exit time $\tau$ increases as one approaches either of these limits, so there is an optimal $\tau$.
Conclusions—We formulated and analyzed a group-driven voter model (GVM) that accounts for the effects of both polyadic interactions and nonlinear interactions within groups. A larger nonlinearity strength $q$ leads to faster consensus in the GVM than in conventional VMs, which exhibit diffusive dynamics. This acceleration of consensus formation depends on the interplay between the nonlinearity strength $q$ and the group size $s$ of hypergraphs. Through mean-field calculations and Monte-Carlo simulations, we demonstrated that the exit time scales logarithmically with system size and that there is an optimal value $q^{*}$ of the nonlinearity strength $q$ that minimizes the exit time. This optimality emerges from a competition between diffusive dynamics when both $q$ and $s$ are small and a slow drift when both $q$ and $s$ are large. This emergent group effect cannot arise in dyadic networks.
We also apply our analytical approach to several variants of our GVM (see the SM SM ): a simplicial GVM, a GVM without allowing duplicate choices, and a GVM with edge-update dynamics in which we simultaneously update the opinions of all nodes that are attached to a hyperedge. In all of these cases, the exit time scales logarithmically with system size, illustrating the robustness of our main theoretical results SM . To further examine the robustness of our results, it is also important to consider additional phenomena. For example, our analysis did not account for heterogeneities in the degree distribution (where the degree of a node is the number of hyperedges it is in). Our preliminary Monte-Carlo calculations SM illustrate that the degree distribution can influence the exit time. Therefore, it will be useful to generalize our mean-field framework to a degree-based mean-field theory Sood and Redner (2005); Sood et al. (2008); Vazquez and Eguíluz (2008) to study their effects analytically. It will also be useful to extend our GVM to encompass more realistic aspects (see, e.g., Redner (2019)) of opinion dynamics. There have been studies of optimal group and team sizes in social psychology Bond (2005); Wheelan (2009), and further studies of our GVM and its generalizations may yield interesting insights about these phenomena.
Acknowledgements.
Acknowledgments—We thank Sid Redner for useful conversations and suggestions. This work was supported in part by National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) [No. 2020R1A2C2003669 (K-IG) and No. 2020R1I1A3068803 (BM)], by a KIAS Individual Grant [No. CG079902] at Korea Institute for Advanced Study (D-SL), by the National Science Foundation (grant number 1922952) through their program on Algorithms for Threat Detection (MAP), and by Agencia Estatal de Investigación (AEI, MCI, Spain) MCIN/AEI/10.13039/501100011033 and Fondo Europeo de Desarrollo Regional (FEDER, UE) under Projects APASOS [PID2021-122256NB-C21] and the Maria de Maeztu Program for Units of Excellence in R&D grant [CEX2021-001164-M] (MSM).